OpenStudy (anonymous):

A rectangular prism has a length x + 2, width 2, height 4, and volume 24. What is the length

OpenStudy (demandandsupply):

volume = area of front surface ie 2x4 times the length

OpenStudy (demandandsupply):

so 2x4=8, then 8 times (x+2)

OpenStudy (anonymous):

macro econ breh?

OpenStudy (demandandsupply):

put that equal to 24 and solve for x, once you've done that, put x in the (x+2) to get the whole length

OpenStudy (anonymous):

Volume is length x width x height. We can figure out the length by dividing the volume by the width and height. 24/4 =6 6/2=3 Length=x+2 3-2=1 x = 1 length=3

OpenStudy (demandandsupply):

@ThickNjuicy ?

OpenStudy (anonymous):

breh chu in macroeconomics dwg?

OpenStudy (demandandsupply):

I study all types of economics, including macro, yes

OpenStudy (anonymous):

tell me bout it dwg